OpenStudy (anonymous):

Which of the following systems is equivalent to the given system? 2/3x - y = 2 x + 1/2y = -3 2x - 3y = 6 and 2x + y = 6 6x - 3y = 6 and 2x + y = 10 2x - 3y = 6 and 2x + y = -6

OpenStudy (advert):

2/3x - y = 2 multiply by 3 2x-3y=6 x + 1/2y = -3 multiply by 2 2x+y=-6 2x-3y=6 2x+y=-6 The third system is the same 2x - 3y = 6 and 2x + y = 6 6x - 3y = 6 and 2x + y = 10 2x - 3y = 6 and 2x + y = -6

OpenStudy (mathmale):

Hints: You are given 2 equations. Multiply all terms of the first equation by 3 to eliminate the fraction (1/3). Multiply all terms of the second equation by 2 to eliminate the fraction (1/2). Note: It's important that you enclose those fractions in parentheses, as I have.

OpenStudy (advert):

There

OpenStudy (advert):

hope i helped